# Flexibility Pricing in Distribution Systems: A Direct Method Aligned with Flexibility Models

### v1

@@ IV-A Modeling of the DSO Flexibility Market Clearing
where ${C_{{\text{Energy}}}}$ is the total energy cost of the distribution system, ${R_{\text{Capacity}}}$ is the revenue from providing reserve capacities, and ${C_{{\text{Flexibility}}}}$ is the total flexbility cost; ${\mathbf{c}}_{\text{energy}}\in\mathbb{R}^{T}$, ${\mathbf{c}}_{\text{ru}}\in\mathbb{R}^{T}$, and ${\mathbf{c}}_{\text{rd}}\in\mathbb{R}^{T}$ denote the vectors composed of the energy price, up-reserve price and down-reserve price, respectively, at the transmission level over the time horizon $[T]$ (assuming that the DSO acts as a price taker at the transmission system, ${\mathbf{c}}_{\text{energy}}$ , ${\mathbf{c}}_{\text{ru}}$, and ${\mathbf{c}}_{\text{rd}}$ are input parameters in the DSO flexibility market); ${{\mathbf{P}}^{\text{ref}}_{0}}\in\mathbb{R}^{T}$ represents the reference power profile at the transmission-distribution interface, i.e., the root node of the distribution system, while ${{\mathbf{R}}^{\text{up}}}\in\mathbb{R}^{T}_{+}$ and ${{\mathbf{R}}^{\text{dn}}}\in\mathbb{R}^{T}_{+}$ represent the up- and down-reserve capacities offered to the transmission system, respectively; aggregators in the distribution system are indexed by $h$, with the set of all aggregators denoted as $\mathcal{A}$; $\Delta\hat{\boldsymbol{\phi}}_{h}$ and $\Delta\check{\boldsymbol{\phi}}_{h}$ denote the activated upward and downward adjustment range of the flexibility trajectory of aggregator $h$, respectively, and $\hat{\mathbf{c}}_{h}$ and $\check{\mathbf{c}}_{h}$ denote the corresponding flexibility cost coefficients, as defined in Section III.
The DSO’s reserve capacity offered to the transmission system must ensure that any adjustment signal within this capacity is feasible with respect to both the distribution network constraints and aggregators’ flexibility constraints, which means the following condition should be satisfied:
@@
where $\mathbb{X}$ represents the feasible region of the power profile ${{\mathbf{P}}_{0}}\in\mathbb{R}^{T}$ at the root node of the distribution system, determined by the distribution power flow constraints and the flexibility models of all aggregators. The upper and lower bounds of ${{\mathbf{P}}_{0}}$ are ${\mathbf{P}}_{0}^{{\text{ref}}}+{{\mathbf{R}}^{{\text{dn}}}}$ and ${\mathbf{P}}_{0}^{{\text{ref}}}-{{\mathbf{R}}^{{\text{up}}}}$, respectively. Down-reserve is related to the plus sign, and up-reserve is related to the minus sign because reserves are defined from the perspective of generators, whereas ${\mathbf{P}}_{0}^{{\text{ref}}}$ and ${{\mathbf{P}}_{0}}$ represent loads with an opposite direction. As we employ the LinDistFlow model to describe the distribution power flow constraints, and the aggregated flexibility model (7) is also linear, the feasible region $\mathbb{X}$ is a convex polytope. Consequently, the condition (13) becomes equivalent to ${\mathbf{P}}_{0}^{{\text{ref}}}-{{\mathbf{R}}^{{\text{up}}}}\in\mathbb{X}$ and ${\mathbf{P}}_{0}^{{\text{ref}}}+{{\mathbf{R}}^{{\text{dn}}}}\in\mathbb{X}$. For the sake of brevity, we let
@@
Constraints in the DSO flexibility market include:
@@
In summary, the DSO flexibility market clearing model is formulated as a linear program: (12), (14)-(28). Solving this model yields the reference power profile ${{\mathbf{P}}^{\text{ref}}_{0}}$ and reserve capacities ${{\mathbf{R}}^{\text{up}}}$ and ${{\mathbf{R}}^{\text{dn}}}$ that the DSO offers to the transmission system, as well as the activated flexibility $\Delta\hat{\boldsymbol{\phi}}_{h}$ and $\Delta\check{\boldsymbol{\phi}}_{h}$ of each aggregator within the distribution system.
IV-B Marginal Flexibility Prices for Settlement
The subsequent task of the flexibility market is the settlement between the DSO and aggregators, which is conducted using marginal pricing. Similar to the concept of Distribution Locational Marginal Prices (DLMPs), we define the sum of the dual variables $\hat{\boldsymbol{\lambda}}_{h}^{\omega}$ and $\check{\boldsymbol{\lambda}}_{h}^{\omega}$ associated with constraints (24) and (25) over $\mathcal{W}$ as the Marginal Flexibility Prices (MFPs):
$\hat{\boldsymbol{\lambda}}_{h}\triangleq\hat{\boldsymbol{\lambda}}_{h}^{\text{ru}}+\hat{\boldsymbol{\lambda}}_{h}^{\text{rd}}\;\;\;\;\text{and}\;\;\;\;\check{\boldsymbol{\lambda}}_{h}\triangleq\check{\boldsymbol{\lambda}}_{h}^{\text{ru}}+\check{\boldsymbol{\lambda}}_{h}^{\text{rd}},$
which reveal the marginal effect of the activated flexibility of aggregator $h$ in reducing the distribution system’s total cost.
When the DSO flexibility market clearing model (12), (14)-(28) is solved, the revenues obtained from the transmission level are distributed to each aggregator based on MFPs. The payment from the DSO to aggregator $h$ is:
@@ IV-B Marginal Flexibility Prices for Settlement
The role of MFPs in the flexibility market is analogous to that of the DLMP in the distribution system’s retail electricity market. First, assuming that the DSO flexibility market is fully competitive, using the MFP for settlement, as in (29), ensures that aggregators’ optimal bidding strategy is to bid their true costs (incentive compatibility). Second, letting $C_{\text{Base}}$ denote the energy cost when the DSO operates on the baseline power profile, we have the following Proposition:
@@
When the DSO flexibility market clearing model (12), (14)-(28) is solved, the following inequality holds:
@@
and equality holds if all voltage constraints in (20) are inactive.
@@
Proposition 2 ensures that, using the MFP for settlement, the total revenue earned from the transmission level is adequate to be distributed to each aggregator (revenue adequacy). In the absence of binding voltage constraints, revenues from energy arbitrage and offering reserve capacities in the transmission system equals the total payments to all aggregators in the distribution system, indicating that the DSO does not make any profits in the flexibility market. Besides, when voltages reach their limits, the DSO will receive a surplus, which is conceptually similar to the TSO’s congestion surplus and can be distributed to market participants or kept by the DSO for operating the flexibility market, upgrading the meters, etc. [4].
Finally, we emphasize that although MFPs are similar to DLMPs in many aspects, MFPs are defined based on the dual variables of the flexibility constraints (24) and (25). This endows MFPs with the following significant property:
Remark 1
The MFPs of different aggregators may differ, even if they are at the same node in the distribution system.
Intuitively, MFPs have this property because they are influenced not only by network constraints but also by the shape of the flexibility region, as defined in (26) and (27). The marginal contributions of different shapes of flexibility regions to the total cost of the distribution system are different. In references [4] and [5], reserves are defined for each aggregator within their flexibility region and then traded in the DSO market. This results in the marginal prices for different aggregators at the same node being identical, all equal to the DLMPs. The theoretical reason for this fact is explained as follows: defining reserves for each aggregator implicitly calculates an inner approximated power range in each time slot (time-decoupled) within the original flexibility region (time-coupled). The feasible region determined by the power range at each time slot is Minkowski additive, meaning that the power range at each time slot can also determine the feasible region at a node, and this range equals the summed ranges of all aggregators at that node. As a result, regardless of how different their original feasible regions are, different aggregators at the same node have the same effect on the nodal power range, thereby contributing identically to the distribution system’s cost and sharing the same marginal price. On the other hand, since we apply the aggregators’ original feasible regions to the market clearing without defining a reserve, those feasible regions are time-coupled and do not have Minkowski additivity. Therefore, changes in the bound parameters of different aggregators contribute differently to the nodal feasible power region, leading to different contributions to the whole system and resulting in different marginal prices.
In contrast, we do not define reserves for each aggregator due to its two disadvantages: first, as discussed in detail in Section 2, it is difficult for aggregators to estimate their reserve cost because using power adjustment ranges only is insufficient for quantifying the flexibility cost of individual DERs; second, defining reserves for each aggregator in the distribution system is unnecessary for participation in the transmission system operation and will lead to conservativeness due to the lack of coordination at the distribution level. Therefore, it is more reasonable to directly define the MFP based on the aggregator’s flexibility constraints (24) and (25). Although this definition leads to the counterintuitive property of MPFs highlighted in Remark 1, MFPs can directly quantify the marginal effect of the aggregator’s time-coupled flexibility on the DSO’s total cost. Therefore, the proposed MFPs ensure a rational settlement of the DSO flexibility market.
@@ V-A Simulation Setup and Flexibility Cost Coefficients
Numerical simulations are carried out on a modified IEEE-33 node system [21] with $T=24$ and $\Delta_{T}=1$ hour. An aggregator, equipped with 20 EVs, 40 HPs, and 1 BESS, is located at each node in the distribution system, except for the root node. Parameters of EVs and BESS are taken from [15], while the data of HPs and corresponding buildings come from Zürich, Switzerland. The aggregated flexibility model (7) is specified as an inner approximated power-energy boundary model [21], ensuring that the aggregated power profile can be feasibly disaggregated to each DER.
@@
In the following subsections, we solve the DSO flexibility market clearing model and analyze the activated flexibility of aggregators, the market settlement results, and the impact of the aggregator’s flexibility cost on the clearing results. We do not include a numerical comparison between the proposed market mechanism and that based on power adjustment ranges because the advantage of the proposed mechanism has been comprehensively discussed in Section II: power adjustment ranges do not adequately quantify the potential inconvenience caused to DER users by flexibility activation, thus failing to be used to formulate the corresponding flexibility costs.
V-B Results of the DSO Flexibility Market Clearing
In the flexibility market clearing, the input data of energy and reserve prices at the transmission level are real data from the Netherlands on Jan 2, 2024, where the energy price varied hourly with an average of 63.06 EUR/MWh, and the up- and down-reserve prices were constant and equal to 12.86 EUR/MW and 14.37 EUR/MW, respectively.
After solving the DSO flexibility market clearing model, the reference power profile and the power profiles corresponding to up- and down-reserves that the DSO offers to the transmission system are shown in Fig. 6. The reference power profile $\mathbf{P}_{0}^{\text{ref}}$ exhibits both upward and downward adjustments from the baseline $\mathbf{P}_{0}^{\text{base}}$, indicating that the DSO shifts the load for energy arbitrage. The gap between $\mathbf{P}_{0}^{\text{ru}}$ and $\mathbf{P}_{0}^{\text{rd}}$ represents reserve capacities. To be more precise, it is the down-reserve (i.e., increasing load), since $\mathbf{P}_{0}^{\text{ru}}$ coincides exactly with $\mathbf{P}_{0}^{\text{ref}}$. This coincidence can be explained as follows. Given $\mathbf{P}_{0}^{\text{ref}}=\mathbf{P}_{0}^{\text{ru}}$, then changing $\mathbf{P}_{0}^{\text{ref}}$ to be higher than $\mathbf{P}_{0}^{\text{ru}}$ (if feasible) will make 1) the flexibility cost remain unchanged, 2) the energy cost increase, 3) the revenue from down-reserve decrease, and 4) the revenue from up-reserve increase. Hence, changing $\mathbf{P}_{0}^{\text{ref}}$ to be higher than $\mathbf{P}_{0}^{\text{ru}}$ would not yield a better solution unless the increase in the revenue from up-reserve exceeds the decrease in the revenue from down-reserve plus the increase in energy cost. This condition, however, is not met in the tested case because the energy prices are significantly higher than the up-reserve prices.
@@ V-B Results of the DSO Flexibility Market Clearing
We take the aggregator at node 2 as an example to illustrate the activated flexibility and the corresponding MFPs (see Fig. 7). The pattern of flexibility activation in Fig. 7, i.e., times and directions of the adjustments, is consistent with the adjustment pattern of power profiles at the root node from the baseline in Fig. 6. For instance, at 20:00, 22:00, and 23:00, $\mathbf{P}_{0}^{\text{ru}}$ and $\mathbf{P}_{0}^{\text{rd}}$ in Fig. 6 exhibit different adjustment directions relative to the baseline, which is reflected in Fig. 7(a), where both the upward and downward power adjustments at these times are nonzero, also correlating with the nonzero energy adjustments in Fig. 7(b) after 20:00. MFPs in Figs. 7(c) and 7(d) show the marginal contribution of power and energy adjustments to the DSO’s total revenue. While the power parts in the aggregator’s flexibility cost coefficients are all zero, the power part in MFP is nonzero because power adjustments contribute to the DSO’s total revenue.
@@
We next discuss the settlement results based on MFPs. First, we study the case without taking the voltage constraints into account, i.e. removing (20) from the problem formulation. In this case, the calculated total revenue of the DSO from energy arbitrage and reserve capacity provision in the transmission system, i.e., the left-hand side of inequality (30), is 4410.98 EUR. Distributing the revenues to each aggregator according to (29), the sum of each aggregator’s payment $R_{h}$, i.e., the right-hand side of inequality (30), is also 4410.98 EUR, exactly the same as the total revenue of the DSO. This result confirms that the proposed MFPs for settlement ensure the DSO’s balance of payments when the voltage constraints are not binding (equivalent to being absent). Then, by adding back the voltage constraints and studying a case where one of these constraints becomes binding, the DSO’s total revenue and the aggregators’ total payments are 4410.98 EUR and 4406.93 EUR, respectively, generating a surplus of 4.05 EUR. The surplus has been discussed right after Proposition 2.
V-C Impact of the Aggregators’ Flexibility Costs on the Market Clearing Results
This section analyzes the impact of the aggregators’ flexibility costs on the clearing results of the DSO’s flexibility market. First, the case where the flexibility costs of all aggregators change is tested. We multiply the cost coefficients of all aggregators by a parameter $\beta$ and analyze how the calculated total net cost and flexibility cost of the DSO change with $\beta$. The results, as shown in Fig. 8, indicate that as the flexibility cost coefficients increase, i.e., $\beta$ increases, the DSO’s total net cost also increases as expected. Second, the DSO’s overall flexibility cost initially increases and then decreases as the flexibility cost coefficients increase. This result is because the total flexibility cost is the product of the cost coefficients and the activated flexibility, and the latter decreases as the flexibility cost coefficients increase.

### v4

@@ IV-A The DSO Optimization Model for Flexibility Activation
where ${C_{{\text{Energy}}}}$ is the total energy cost of the distribution system, ${R_{\text{Capacity}}}$ is the revenue from providing reserve capacities, and ${C_{{\text{Flexibility}}}}$ is the total flexbility cost; ${\mathbf{c}}_{\text{energy}}\in\mathbb{R}^{T}$, ${\mathbf{c}}_{\text{ru}}\in\mathbb{R}^{T}$, and ${\mathbf{c}}_{\text{rd}}\in\mathbb{R}^{T}$ denote the vectors composed of the energy price, up-reserve price and down-reserve price, respectively, at the transmission level over the time horizon $[T]$ (assuming that the DSO acts as a price taker at the transmission system, ${\mathbf{c}}_{\text{energy}}$ , ${\mathbf{c}}_{\text{ru}}$, and ${\mathbf{c}}_{\text{rd}}$ are input parameters at the DSO level); ${{\mathbf{P}}^{\text{ref}}_{0}}\in\mathbb{R}^{T}$ represents the reference power profile at the TSO-DSO interface, i.e., the root node of the distribution system, while ${{\mathbf{R}}^{\text{up}}}\in\mathbb{R}^{T}_{+}$ and ${{\mathbf{R}}^{\text{dn}}}\in\mathbb{R}^{T}_{+}$ represent the up- and down-reserve capacities offered to the TSO, respectively; aggregators in the distribution system are indexed by $h$, with the set of all aggregators denoted as $\mathcal{A}$; $\Delta\hat{\boldsymbol{\phi}}_{h}$ and $\Delta\check{\boldsymbol{\phi}}_{h}$ denote the activated upward and downward adjustment range of the flexibility trajectory of aggregator $h$, respectively, and $\hat{\mathbf{c}}_{h}$ and $\check{\mathbf{c}}_{h}$ denote the corresponding flexibility cost coefficients, as defined in Section III.
The DSO’s reserve capacity offered to the TSO must ensure that any adjustment signal within this capacity is feasible with respect to both the distribution network constraints and aggregators’ flexibility constraints, which means the following condition should be satisfied:
@@
where $\mathbb{X}$ represents the feasible region of the power profile ${{\mathbf{P}}_{0}}\in\mathbb{R}^{T}$ at the root node of the distribution system, determined by the distribution power flow constraints and the flexibility models of all aggregators. The upper and lower bounds of ${{\mathbf{P}}_{0}}$ are ${\mathbf{P}}_{0}^{{\text{ref}}}+{{\mathbf{R}}^{{\text{dn}}}}$ and ${\mathbf{P}}_{0}^{{\text{ref}}}-{{\mathbf{R}}^{{\text{up}}}}$, respectively. Down-reserve is related to the plus sign, and up-reserve is related to the minus sign because reserves are defined from the perspective of generators, whereas ${\mathbf{P}}_{0}^{{\text{ref}}}$ and ${{\mathbf{P}}_{0}}$ represent loads with an opposite direction. As we employ the LinDistFlow model to describe the distribution power flow constraints, the closed form of $\mathbb{X}$ can also be expressed as (9), but the parameters $\overline{\boldsymbol{\phi}}$ and $\underline{\boldsymbol{\phi}}$ will be affected by the distribution network constraints (See our previous work [24] for a detailed theoretical analysis). Since all the components in $\mathbf{U}$ are no less than zero, ${\mathbf{P}}_{0}^{{\text{ref}}}-{{\mathbf{R}}^{{\text{up}}}}\in\mathbb{X}$ and ${\mathbf{P}}_{0}^{{\text{ref}}}+{{\mathbf{R}}^{{\text{dn}}}}\in\mathbb{X}$ is sufficient to ensure that any ${{\mathbf{P}}_{0}}$ such that ${\mathbf{P}}_{0}^{{\text{ref}}}-{{\mathbf{R}}^{{\text{up}}}}\leq{{\mathbf{P}}_{0}}\leq{\mathbf{P}}_{0}^{{\text{ref}}}+{{\mathbf{R}}^{{\text{dn}}}}$ also belongs to $\mathbb{X}$. Consequently, the condition (15) becomes equivalent to ${\mathbf{P}}_{0}^{{\text{ref}}}-{{\mathbf{R}}^{{\text{up}}}}\in\mathbb{X}$ and ${\mathbf{P}}_{0}^{{\text{ref}}}+{{\mathbf{R}}^{{\text{dn}}}}\in\mathbb{X}$. For the sake of brevity, we let
@@
Constraints in the distribution system include:
@@
In summary, the DSO optimization model for flexibility activation is formulated as a linear programming: (14), (16)-(30). Solving this model yields the reference power profile ${{\mathbf{P}}^{\text{ref}}_{0}}$ and reserve capacities ${{\mathbf{R}}^{\text{up}}}$ and ${{\mathbf{R}}^{\text{dn}}}$ that the DSO offers to the TSO, as well as the activated flexibility $\Delta\hat{\boldsymbol{\phi}}_{h}$ and $\Delta\check{\boldsymbol{\phi}}_{h}$ of each aggregator within the distribution system.
IV-B Marginal Flexibility Prices for Revenue Allocation
Next, the revenues earned from the TSO should be allocated to each aggregator in the distribution system. A reasonable method is to allocate the revenues based on the marginal contributions of each aggregator’s activated flexibility to the total revenue of the distribution system, which encourages each aggregator to report its true costs (incentive compatibility). We propose the concept of Marginal Flexibility Prices (MFPs) to quantify the marginal contributions of each aggregator’s activated flexibility to the total revenue of the DSO (the value of flexibility). The MFP of aggregator $h$ is defined as the sum of the dual variables $\hat{\boldsymbol{\lambda}}_{h}^{\omega}$ and $\check{\boldsymbol{\lambda}}_{h}^{\omega}$ associated with constraints (26) and (27) over $\mathcal{W}$:
$\hat{\boldsymbol{\lambda}}_{h}\triangleq\hat{\boldsymbol{\lambda}}_{h}^{\text{ru}}+\hat{\boldsymbol{\lambda}}_{h}^{\text{rd}}\;\;\;\;\text{and}\;\;\;\;\check{\boldsymbol{\lambda}}_{h}\triangleq\check{\boldsymbol{\lambda}}_{h}^{\text{ru}}+\check{\boldsymbol{\lambda}}_{h}^{\text{rd}},$ (31)
which reveal the marginal effect of the activated flexibility of aggregator $h$ in reducing the distribution system’s net cost. Allocating the total revenue of the DSO according to the MFPs, aggregator $h$ can obtain a payment of:
@@ IV-B Marginal Flexibility Prices for Revenue Allocation
The role of MFPs in the flexibility activation process is analogous to that of the distribution locational marginal prices (DLMPs) in the distribution system’s retail electricity market. In addition to incentive compatibility, allocating the revenues based on MFPs also has the following property:
@@
The solution to the DSO optimization model (14), (16)-(30) makes the following inequality hold:
@@
where, $C_{\text{Base}}$ denote the energy cost when the DSO operates on the baseline power profile. The equality holds if all voltage constraints in (22) are not binding.
@@
Proposition 2 ensures that, using the MFPs for revenue allocation, the revenues earned from the TSO are adequate to be allocated to each aggregator (revenue adequacy). Besides, in the absence of binding voltage constraints, the revenues obtained from the TSO can be fully allocated to all aggregators in the distribution system, indicating that the DSO does not make any profits in the flexibility activation process. When voltages reach their limits, the DSO will receive a surplus, which is conceptually similar to the TSO’s congestion surplus and can be distributed to the aggregators or kept by the DSO for operating the flexibility coordination, upgrading the meters, etc. [3].
@@ V-A Simulation Setup and Flexibility Cost Coefficients
Numerical simulations are carried out on a modified IEEE-33 node system [17] with $T=24$ and $\Delta_{T}=1$ hour. There are 32 aggregators distributed at each node except for the root node in the distribution system, each equipped with 20 EVs, 40 HPs, and 1 BESS. Parameters of EVs and BESS are taken from [16], while the data of HPs and corresponding buildings come from Zürich, Switzerland. The aggregated flexibility model (9) is specified as an inner approximated power-energy boundary model [17] (with $24T-$ 2 constraints per aggregator), ensuring that the aggregated power profile can be feasibly disaggregated to each DER.
@@
In the following subsections, we solve the DSO optimization model and analyze the activated flexibility of aggregators, the revenue allocation results, and the impact of the aggregator’s flexibility cost on the results. We have considered making a numerical comparison between the proposed methodologies and those in existing studies. However, in the simulation of many existing studies, aggregators are assumed to only report their cost for providing power adjustment range at each time slot to the DSO, without explaining where these cost parameters come from nor reflecting the impact of flexibility activation on the DER devices or/and their users [18, 3, 4, 19, 20]. As analyzed in Section II, power adjustment ranges cannot adequately quantify the flexibility costs associated with time-coupled flexibility regions. Therefore, the DSO optimization models and revenue allocation strategies presented in these works are fundamentally unsuitable for the scenario where the flexibility cost is quantified in a time-coupled manner, making numerical comparisons unviable. The advantage of the proposed flexibility quantification method over traditional power-range-based approaches lies in its consistency with the flexibility models and capability of reflecting the potential inconvenience caused to DER users offering flexibility, which has been comprehensively discussed in Section II. Thus, the numerical simulations here mainly verify the reasonableness of flexibility activation and revenue allocation results calculated by the proposed DSO optimization model and revenue allocation strategy.
V-B Results of Flexibility Activation and Revenue Allocation
In the DSO optimization model, the input data of energy and reserve prices at the transmission level are real data from the Netherlands on Jan 2, 2024, where the energy price varied hourly with an average of 63.06 EUR/MWh, and the up- and down-reserve prices were constant and equal to 12.86 EUR/MW and 14.37 EUR/MW, respectively.
After solving the DSO optimization model, the DSO’s reference power profile and the power profiles corresponding to up- and down-reserve deployments are shown in Fig. 6. The reference power profile $\mathbf{P}_{0}^{\text{ref}}$ exhibits both upward and downward adjustments from the baseline $\mathbf{P}_{0}^{\text{base}}$, indicating that the DSO shifts the load for energy arbitrage. The gap between $\mathbf{P}_{0}^{\text{ru}}$ and $\mathbf{P}_{0}^{\text{rd}}$ represents reserve capacities. To be more precise, it is the down-reserve (i.e., increasing load), since $\mathbf{P}_{0}^{\text{ru}}$ coincides exactly with $\mathbf{P}_{0}^{\text{ref}}$. This coincidence can be explained as follows. Given $\mathbf{P}_{0}^{\text{ref}}=\mathbf{P}_{0}^{\text{ru}}$, then changing $\mathbf{P}_{0}^{\text{ref}}$ to be higher than $\mathbf{P}_{0}^{\text{ru}}$ (if feasible) will make 1) the flexibility cost remain unchanged, 2) the energy cost increase, 3) the revenue from down-reserve decrease, and 4) the revenue from up-reserve increase. Hence, changing $\mathbf{P}_{0}^{\text{ref}}$ to be higher than $\mathbf{P}_{0}^{\text{ru}}$ would not yield a better solution unless the increase in the revenue from up-reserve exceeds the decrease in the revenue from down-reserve plus the increase in energy cost. This condition, however, is not met in the tested case because the energy prices are significantly higher than the up-reserve prices.
@@ V-B Results of Flexibility Activation and Revenue Allocation
We take the aggregator at node 2 as an example to illustrate the activated flexibility and the corresponding MFPs (see Fig. 7). The pattern of flexibility activation in Fig. 7, i.e., times and directions of the adjustments, is consistent with the adjustment pattern of power profiles at the root node from the baseline in Fig. 6. For instance, at 20:00, 22:00, and 23:00, $\mathbf{P}_{0}^{\text{ru}}$ and $\mathbf{P}_{0}^{\text{rd}}$ in Fig. 6 exhibit different adjustment directions relative to the baseline, which is reflected in Fig. 7(a), where both the upward and downward power adjustments at these times are nonzero, also correlating with the nonzero energy adjustments in Fig. 7(b) after 20:00. The power and energy parts of the aggregator’s MFPs are shown in Figs. 7(c) and 7(d). While the power part in the aggregator’s flexibility cost coefficients are all zero, the power part in MFPs is nonzero because power adjustments contribute to the DSO’s total revenue.
@@
We next discuss the revenue allocation results based on MFPs. First, we study the case without taking the voltage constraints into account, i.e. removing (22) from the problem formulation. In this case, the total revenue of the DSO from energy arbitrage and reserve capacity provision in the transmission system, i.e., the left-hand side of inequality (33), is 4410.98 EUR. Distributing the revenues to each aggregator according to (32), the sum of each aggregator’s payments $R_{h}$, i.e., the right-hand side of inequality (33), is also 4410.98 EUR, exactly the same as the total revenue of the DSO. This result confirms that the proposed MFPs for settlement ensure the DSO’s balance of payments when the voltage constraints are not binding (equivalent to being absent). However, if we use only the dual variables corresponding to power range constraints for revenue allocation, the total payment allocated to all aggregators is only 1118.46 EUR. This result indicates that the primary factor affecting the total revenue of the distribution system is not the power adjustment range but the energy adjustment range in this simulation. Therefore, it is necessary to use MFPs, which include both the power and energy parts, for revenue allocation.
Then, by adding back the voltage constraints and studying a case where one of these constraints becomes binding, the DSO’s total revenue and the aggregators’ total payments are 4410.98 EUR and 4406.93 EUR, respectively, generating a surplus of 4.05 EUR. The surplus has been discussed right after Proposition 2.
V-C Impact of the Aggregators’ Flexibility Costs on the Results
This section analyzes the impact of the aggregators’ flexibility costs on the DSO’s flexibility activation results. First, the case where the flexibility costs of all aggregators change is tested. We multiply the cost coefficients of all aggregators by a parameter $\beta$ and analyze how the calculated total net cost and flexibility cost of the DSO change with $\beta$. The results, as shown in Fig. 8, indicate that as the flexibility cost coefficients increase, i.e., $\beta$ increases, the DSO’s total net cost also increases as expected. Second, the DSO’s overall flexibility cost initially increases and then decreases as the flexibility cost coefficients increase. This result is because the total flexibility cost is the product of the cost coefficients and the activated flexibility, and the latter decreases as the flexibility cost coefficients increase.